# Casimir Arc–Plate Geometry: Computational Analysis of Thickness Constraints for Gold and Silver Nanomembranes in MEMS Applications

### v1

@@ Introduction
In 1947, in “The Influence of Retardation on the London-van der Waals Forces,” H. B. G. Casimir and D. Polder observed the London-van der Waals interaction of two neutral atoms. Their initial approach involved the use of second-order perturbation theory, considering the interaction of a neutral atom with a perfectly conducting wall, to measure the van der Waals force at increased separations [1]. Further, Casimir treated the same problem in terms of vacuum; he calculated the zero-point energy of the electromagnetic field in the presence of molecules and compared it to the corresponding vacuum energy. Finally, to simplify this setup, Casimir reviewed the same interaction between two conductive plates rather than between molecules [2]. This observation led directly to the prediction of the Casimir effect.
@@ 1.1 The Casimir Effect
The Casimir effect is observed in the attractive force between two perfectly conducting, uncharged parallel plates, measured as in equation 1 for distance $d$. The speed of light, $c$, transforms the electromagnetic mode wavelength into a frequency, which is further converted to energy by the reduced Planck constant. The negative sign represents the attractive force (later we will observe the conditions under which the Casimir force can act as a repulsive force) [3].
@@
This effect is a consequence of the quantum vacuum: the vacuum fluctuations in the space between the plate and the outside space cause a pressure difference. In his calculations, Casimir proved that even at any finite volume, the zero-point energy of an electromagnetic vacuum is infinite, and changes in the energy can be identified. Specifically, this observation estimates the potential energy required to reduce the separation between the two plates from $d\rightarrow\infty$ , to distance $d$. This derivation fundamentally corresponds to the Casimir force [4]. The complete derivation can be reviewed from numerous sources [3], [4].
@@ 1.2 Geometry
Although the derivation of the Casimir effect describes the interaction between two parallel plates, a variety of different geometries have arisen in the experimental setups. Configurations, such as plate-to-sphere [5], sphere-to-sphere [6], and plate-to-cylinder [7], have been experimentally evaluated. Owing to the challenge of keeping the geometry of the two plates parallel, many of the later performed experiments experienced a change in geometry, namely, the sphere-to-plate geometry. However, the vacuum energy estimate is geometry-dependent. Thus, the variation in the geometry of one of the two surfaces modifies the electromagnetic modes, resulting in varied energy differences. To express the Casimir effect for a curved surface, the proximity force approximation (PFA) method approximates the surface as a set of infinitesimal parallel plates, calculates the Casimir force for each pair of plates at the respective distance, and integrates the contributions of all pairs to obtain the total force. For separation $d$ and curvature $R$, $d<<R$ must be true [8]. The simplified formula for the sphere-to-plate geometry, acquired from the PFA method, is described as in equation 2, where $R$ is the radius of the sphere and the condition $d<<R$ is met.
@@
Theoretical analysis involves the geometries of a cylinder to a plate, parabolic cylinder to a plate, paraboloid to a plate, and sinusoidal corrugation to a plate [8]. However, PFA treats the locally flat regions as parallel to the opposing surface, and as a result it is considered an uncontrolled approximation. Next to leading order (NTLO) corrections are adapted to consider the effects of a curved geometry: a derivative expansion of the geometrical profile of the surface is considered up to the second term of the expansion [8]. Moreover, the study by [9] advanced corrections for finite temperature, two smooth surfaces, and various other cases.
@@ 1.3 MEMS
Owing to the microscale and magnitude of force compared to the scale, the Casimir effect has been applied in micro-electro-mechanical systems (MEMS) and sensors. The Casimir effect serves in quantum metrology, the approach of measuring parameters with increased accuracy using quantum phenomena. Quantum metrology has important applications in the detection of weak magnetic fields. The intended use of the Casimir effect in such devices, specifically magnetic sensors, utilized to assess brain and cardiac activity, cancerous tumors, and cosmic monitoring, is discussed in [10]. In these systems, the Casimir effect augments the measurement resolution. Moreover, the study by [11] effectively embedded a Casimir geometry  [^1] Reference [11] theoretically calculates the Casimir force, and further experimentally verifies the results. We would like to note that in this study, the arc geometry would make use of the PFA with NTLO due to its geometry properties. Specifically, the slope of the sphere is constant, while the slope of an arc can vary more rapidly where local distances change nonlinearly. on a commercial MEMS inertial sensor and measured the Casimir force. Monitoring the Casimir effect using a commercial MEMS device suggests its implementation as an engineering instrument in commercial MEMS. Furthermore, [12] modeled a quantum displacement amplifier that converts DC to a much larger amplitude and frequency change in the AC. The Casimir effect is used to extract energy from the vacuum and supply it to a mechanical parametric amplifier. Combining the quantum-mechanical Casimir effect and MEMS results in a quantum metrology device, which can be applied to obtain zeptometer-level precision measurements.
@@
However, the major challenge in the application of the Casimir force in MEMS is stiction, referred to as Casimir pull-in, which can lead to device failure. In [10], the authors not only amplified the signals up to 10,000 times using the Casimir effect, but also prevented Casimir pull-in through time-delay-based parametric amplification. This is achieved by modulating the MEMS magnet oscillation parameter at specific frequencies and intentionally timing the oscillations to avoid collapse while achieving amplification. This displacement modulation counteracts the Casimir force, thereby ensuring that the components do not collapse.
@@
While prior theoretical studies have focused on plate-to-sphere or plate-to-cylinder configurations, arc-to-plate geometry, particularly for gold and silver nanomembranes, remains unexplored despite its practical relevance for MEMS applications. This study introduces a computational method to determine the maximum nanomembrane thickness that allows the arc to undergo curvature reversal under the Casimir force, enabling new insights for MEMS actuation. Additionally, these calculations provide a lower-bound constraint that can be applied in microfabrication to prevent stiction, demonstrating its practical use in improving the reliability of MEMS.
This study focuses on analytical modeling of Casimir-induced constraints in curved nanomembrane geometries. While experimental validation and finite-element simulations would be valuable for quantitative device optimization, the present study aims to establish conservative theoretical bounds relevant to ultra-thin membranes. We identify numerical validation of the Casimir forces as an important direction for future work.
@@ 2.1.1 The Quantum Vacuum
Casimir recognized quantum electrodynamics (QED), where a vacuum has a lower energy limit for the system to be conserved, known as zero-point energy. Zero-point energy is essentially the sum of all possible frequencies, which explains its infinite magnitude. This is expressed in Peter W. Milonni’s “The Quantum Vacuum: An Introduction to Quantum Electrodynamics”, where the energy $\sum_{k\lambda}\frac{1}{2}\hbar w_{k}$ is infinite [4].
[…]
@@
To maintain the energy-time uncertainty relation, quantum vacuum fluctuations arise simply as random, temporary changes in energy. The appearance and disappearance of virtual particles serve as mathematical representations of disturbances in the field. As a result of the vacuum field, many phenomena can be observed, some of which are the Lamb shift, atoms spontaneously emitting photons, and Casimir force [4].
@@ 2.1.2 Boundary Conditions
In “On the attraction between two perfectly conducting plates”(1948) Casimir shows that the electromagnetic field experiences changes to satisfy the boundary conditions introduced by the plates [14]. Specifically, a pressure difference was observed, which is the cause of the assumption that the plates have infinite conductivity. The perfect conductivity of the plates creates an electromagnetic field within the conductor, $E=0$. To satisfy Maxwell’s boundary conditions, the tangential component vanishes [15]. Mathematically, this corresponds to the Dirichlet boundary condition, which requires a field to be fixed at zero in the plates. Moreover, the enclosed space between the plates allows only wavelengths of certain $\lambda$ to propagate in the field. For $\lambda$, equation 5 expresses the boundary conditions that should be satisfied, where, $\lambda$ is the allowed wavelength of the field between the plates, $d$ is the distance between the plates, and $n$ is a positive integer ($n=1,2,3,$...) representing the mode number.
@@ 2.1.3 Repulsive Casimir effect
Although the Casimir force is attractive at the origin, a repulsive Casimir force was measured in [16] when solid materials were immersed in a fluid. Munday, Capasso and Parsegian considered the dielectric response functions, $\epsilon$, of the materials. They observed that in a configuration of $\epsilon_{1}>\epsilon_{3}>\epsilon_{2}$ , where $\epsilon_{3}$ represents the medium and $\epsilon_{1}$ and $\epsilon_{2}$ the materials, the force is repulsive. These findings suggest applications such as quantum levitation, allowing the development of ultra-low friction devices and sensors [16]. In addition, the repulsive Casimir force was observed in [17] in a system consisting of a gold surface and a Teflon coated gold surface, immersed in ethanol.
Other methods for obtaining a repulsive Casimir force involve modifications of the geometry, such as the proposed glide-symmetric geometry [18]. We observed that a change in geometry altered the allowed field modes. The repulsive Casimir force case is usually achieved by introducing an asymmetry between the two surfaces. Thus, both attractive and repulsive Casimir forces can be applied to MEMS.
@@ 2.1.4 Experiments and Materials
With a conductivity value of $4.5\times 10^{7}S/m$, gold is the most commonly used material because of its ability to be modified to achieve minimal roughness. Methods involve template stripping as suggested in [19], achieving a root mean square roughness of less than $0.4$ nm, which is crucial for the accuracy of measurements at nanoscale distances. Moreover, [20] distinguished the difference in the observed force when both bodies were made of Au, compared to when one body was made of Au and the other of Si. The force between the Au sphere and Si plate geometry with respect to the Au sphere and Au plate configuration was represented in a ratio of $0.74$ at separations of $100$ nm and $0.63$ at separations of $200$ nm , as recognized in [21].
Silver has an even higher conductivity value of $6.3\times 10^{7}S/m$, which should theoretically increase the magnitude of the Casimir force. However, owing to overtime oxidation, the surface roughness increases and, at the nanoscale, this will affect the accuracy of measurements, making gold preferred in experimental setups. The oxidation rate of silver increases with a decrease in the size of silver particles, notably for diameters $d_{p}\leq 6$ nm [22]. Furthermore, the oxidation process is shape-dependent, as described in [23], where the oxidation process of silver nanostructures is observed, specifically for spheres, cubes, disks, cylinders, triangles, and pyramids. The nanospheres had the lowest level of oxidation among all the geometries tested [23]. Selecting specific silver nanostructures according to their oxidation characteristics can improve the design of MEMS and NEMS. Thus, the material properties affect the accuracy of the theoretical results compared to the experimental evaluation.
The Casimir force is observed as dominant in systems at the microscale and has been further experimentally verified. The Casimir force is observed as dominant in systems at the microscale and has been further experimentally verified. Lamoreaux conducted the first experimental measurement of the Casimir force between a flat plate and a sphere using a torsion pendulum for separations of $0.6$ to $6\mu m$. Both components of the composed geometry were coated with a $0.5\mu m$ Cu layer on each side, and further $0.5\mu m$ Au was evaporated on the faces undergoing attraction. The theoretical results agreed with the experimental measurements within an error of $5\%$ [5]. Another common method involves atomic force microscopy (AFM), owing to the high accuracy of measurements that it allows. U. Mohideen and A. Roy first implemented the AFM method for a plate-to-sphere geometry with a separation range of $0.1$ to $0.9\mu m$, and obtained an accuracy of $1\%$ [24].
@@ 2.2 Theoretical Framework
First, a Casimir geometry consisting of an arc and a plate was introduced. The arc is essentially a finite curved nanomembrane, with length which we approximate to $6\mu m$ (due to the large radius of $R=100\mu m),$ concave to a finite plate with length 6 $\mu m$ (figure 1(a)). Both surfaces were fixed so that the Casimir interaction could not bring them together. The nanomembrane is modeled as a thin elastic arc with clamed endpoints, as a representation of a membrane fixed to a rigid MEMS frame. Clamped boundary conditions are assumed at the ends of the arc, which constrains both displacement and rotation. These assumptions are consistent with typical MEMS design [25]. Future work will extend these results through formal stability analysis. Figure 1(a) represents the arc-plate geometry at a separation of $0.1\mu m$ of the two surfaces, and figure 1(b) shows the arc-plate geometry at a distance $1\mu m$. Reference [26] shows that Casimir edge effects for finite plates are negligible when the finite extent of the plate in the direction parallel to its surface is sufficiently small. In this study, we consider the plate and arc to be sufficiently small in the direction parallel to their surfaces to satisfy this condition, allowing edge effects to be safely neglected throughout our analysis.
@@
The distance range between the arc and the plate, considered from the farthest point of the arc, is described as $d\in[0.1-1]\mu m$. Although the Casimir force increases with decreasing distance, most experiments operate in the micro- or sub-micro range, as in [24], [27]. Decreasing the distance to the nanoscale introduces challenges such as effects due to surface roughness and inability to achieve precision as the Casimir force becomes uncontrollably strong.
@@
The material choice is explained in Section 2.1.4. ”Experiments and Materials” of this paper. The values for the Poisson’s ratio and Young’s modulus were taken into account in the case of a thin film for both gold and silver.

### v2

@@ Introduction
In 1947, in “The Influence of Retardation on the London-van der Waals Forces,” H. B. G. Casimir and D. Polder observed the London-van der Waals interaction of two neutral atoms. Their initial approach involved the use of second-order perturbation theory, considering the interaction of a neutral atom with a perfectly conducting wall, to measure the van der Waals force at increased separations (casimir_influence_1948, ). Further, Casimir treated the same problem in terms of vacuum; he calculated the zero-point energy of the electromagnetic field in the presence of molecules and compared it to the corresponding vacuum energy. Finally, to simplify this setup, Casimir reviewed the same interaction between two conductive plates rather than between molecules (jaffe_casimir_2005, ). This observation led directly to the prediction of the Casimir effect.
@@ 1.1 The Casimir Effect
The Casimir effect is observed in the attractive force between two perfectly conducting, uncharged parallel plates, measured as in equation 1 for distance $d$. The speed of light, $c$, transforms the electromagnetic mode wavelength into a frequency, which is further converted to energy by the reduced Planck constant. The negative sign represents the attractive force (later we will observe the conditions under which the Casimir force can act as a repulsive force) (lamoreaux_casimir_2005, ).
@@
This effect is a consequence of the quantum vacuum: the vacuum fluctuations in the space between the plate and the outside space cause a pressure difference. In his calculations, Casimir proved that even at any finite volume, the zero-point energy of an electromagnetic vacuum is infinite, and changes in the energy can be identified. Specifically, this observation estimates the potential energy required to reduce the separation between the two plates from $d\rightarrow\infty$ , to distance $d$. This derivation fundamentally corresponds to the Casimir force (milonni_quantum_1994, ). The complete derivation can be reviewed from numerous sources (lamoreaux_casimir_2005, ), (milonni_quantum_1994, ).
@@ 1.2 Geometry
Although the derivation of the Casimir effect describes the interaction between two parallel plates, a variety of different geometries have arisen in the experimental setups. Configurations, such as plate-to-sphere (lamoreaux_demonstration_1997, ), sphere-to-sphere (garrett_measurement_2018, ), and plate-to-cylinder (emig_casimir_2006, ), have been experimentally evaluated. Owing to the challenge of keeping the geometry of the two plates parallel, many of the later performed experiments experienced a change in geometry, namely, the sphere-to-plate geometry. However, the vacuum energy estimate is geometry-dependent. Thus, the variation in the geometry of one of the two surfaces modifies the electromagnetic modes, resulting in varied energy differences. To express the Casimir effect for a curved surface, the proximity force approximation (PFA) method approximates the surface as a set of infinitesimal parallel plates, calculates the Casimir force for each pair of plates at the respective distance, and integrates the contributions of all pairs to obtain the total force. For separation $d$ and curvature $R$, $d<<R$ must be true (fosco_proximity_2011, ). The simplified formula for the sphere-to-plate geometry, acquired from the PFA method, is described as in equation 2, where $R$ is the radius of the sphere and the condition $d<<R$ is met.
@@
Theoretical analysis involves the geometries of a cylinder to a plate, parabolic cylinder to a plate, paraboloid to a plate, and sinusoidal corrugation to a plate (fosco_proximity_2011, ). However, PFA treats the locally flat regions as parallel to the opposing surface, and as a result it is considered an uncontrolled approximation. Next to leading order (NTLO) corrections are adapted to consider the effects of a curved geometry: a derivative expansion of the geometrical profile of the surface is considered up to the second term of the expansion (fosco_proximity_2011, ). Moreover, the study by (fosco_casimir_2024, ) advanced corrections for finite temperature, two smooth surfaces, and various other cases.
Unlike the sphere-plate geometry commonly analyzed in literature, including Ref. (stange_building_2019, ), the arc–plate configuration considered in this work exhibits a spatially varying slope and a nonlinear variation of the local separation along the nanomembrane. In a sphere–plate geometry the surface slope changes smoothly and remains nearly constant over the interaction region, whereas for an arc geometry the slope can vary more rapidly, particularly near regions where the nanomembrane comes closest to the plate. As a result, the leading-order PFA alone may not fully capture the geometric dependence of the Casimir interaction. For this reason, the thickness estimates could benefit from NTLO corrections to ensure a consistent description of the Casimir energy in the arc–plate geometry, which is described later in our study.
@@ 1.3 MEMS
Owing to the microscale and magnitude of force compared to the scale, the Casimir effect has been applied in micro-electro-mechanical systems (MEMS) and sensors. The Casimir effect serves in quantum metrology, the approach of measuring parameters with increased accuracy using quantum phenomena. Quantum metrology has important applications in the detection of weak magnetic fields. The intended use of the Casimir effect in such devices, specifically magnetic sensors, utilized to assess brain and cardiac activity, cancerous tumors, and cosmic monitoring, is discussed in (javor_analysis_2021, ). In these systems, the Casimir effect augments the measurement resolution. Moreover, the study by (stange_building_2019, ) effectively embedded a Casimir geometry on a commercial MEMS inertial sensor and measured the Casimir force. Monitoring the Casimir effect using a commercial MEMS device suggests its implementation as an engineering instrument in commercial MEMS. Furthermore, (javor_zeptometer_2022, ) modeled a quantum displacement amplifier that converts DC to a much larger amplitude and frequency change in the AC. The Casimir effect is used to extract energy from the vacuum and supply it to a mechanical parametric amplifier. Combining the quantum-mechanical Casimir effect and MEMS results in a quantum metrology device, which can be applied to obtain zeptometer-level precision measurements.
@@
However, the major challenge in the application of the Casimir force in MEMS is stiction, referred to as Casimir pull-in, which can lead to device failure. In (javor_analysis_2021, ), the authors not only amplified the signals up to 10,000 times using the Casimir effect, but also prevented Casimir pull-in through time-delay-based parametric amplification. This is achieved by modulating the MEMS magnet oscillation parameter at specific frequencies and intentionally timing the oscillations to avoid collapse while achieving amplification. This displacement modulation counteracts the Casimir force, thereby ensuring that the components do not collapse.
@@
While prior theoretical studies have focused on plate-to-sphere or plate-to-cylinder configurations, arc-to-plate geometry, particularly for gold and silver nanomembranes, remains unexplored despite its practical relevance for MEMS applications. This study introduces a theoretical estimate to determine the maximum nanomembrane thickness that allows the arc to undergo curvature reversal under the Casimir force, enabling new insights for MEMS actuation. Additionally, these calculations provide a lower-bound constraint that can be applied in microfabrication to prevent stiction, demonstrating its practical use in improving the reliability of MEMS.
@@ 2.1.1 The Quantum Vacuum
Casimir recognized quantum electrodynamics (QED), where a vacuum has a lower energy limit for the system to be conserved, known as zero-point energy. Zero-point energy is essentially the sum of all possible frequencies, which explains its infinite magnitude. This is expressed in Peter W. Milonni’s “The Quantum Vacuum: An Introduction to Quantum Electrodynamics”, where the energy $\sum_{k\lambda}\frac{1}{2}\hbar w_{k}$ is infinite (milonni_quantum_1994, ).
To understand the concept of zero-point energy, one should recognize Heisenberg’s uncertainty principle, which is a necessary condition that combines three interrelated uncertainty relations for position and momentum (milonni_quantum_1994, ). First, the uncertainty principle constrains the widths of the position and momentum probability distribution in any quantum state; thus, when the position distribution is localized, the momentum distribution must be larger in the width area, and vice versa. Second, a disturbance uncertainty appears: there is a trade-off between measurements of the position and momentum; precisely measuring one causes a disturbance in the other. Third, owing to the non-commutativity of the position and momentum in the Heisenberg uncertainty bound, the constraints imposed by the joint approximation prevent the simultaneous determination of both quantities (busch_heisenbergs_2007, ). All three conditions describe Heisenberg’s uncertainty, where the zero-point energy is explained specifically by the energy-time relation (equation 3), which is a specific form of the generalized Heisenberg inequality (equation 4). To explain vacuum, an inverse scenario is observed: if the system has energy, s.t. $\Delta E=0$ , the time would have to reach infinity. Considering a finite-time system, the energy value remains uncertain. Therefore, energy there must be present in the vacuum. In equation 3 $\Delta E$ is the uncertainty in energy, $\Delta t$ is the uncertainty in time, and $\hbar$ is the reduced Planck constant. In equation 4 $\sigma_{a}$ and $\sigma_{b}$ are the standard deviations (uncertainties) of the observables $A$ and $B$, [A,B]=ABBA[A,B]=AB-BA is their commutator, $\langle[A,B]\rangle$ denotes its expectation value, and $i$ is the imaginary unit.
@@
To maintain the energy-time uncertainty relation, quantum vacuum fluctuations arise simply as random, temporary changes in energy. The appearance and disappearance of virtual particles serve as mathematical representations of disturbances in the field. As a result of the vacuum field, many phenomena can be observed, some of which are the Lamb shift, atoms spontaneously emitting photons, and Casimir force (milonni_quantum_1994, ).
@@ 2.1.2 Boundary Conditions
In “On the attraction between two perfectly conducting plates”(1948) Casimir shows that the electromagnetic field experiences changes to satisfy the boundary conditions introduced by the plates (casimir_attraction_nodate, ). Specifically, a pressure difference was observed, which is the cause of the assumption that the plates have infinite conductivity. The perfect conductivity of the plates creates an electromagnetic field within the conductor, $E=0$. To satisfy Maxwell’s boundary conditions, the tangential component vanishes (jackson_classical_2009, ). Mathematically, this corresponds to the Dirichlet boundary condition, which requires a field to be fixed at zero in the plates. Moreover, the enclosed space between the plates allows only wavelengths of certain $\lambda$ to propagate in the field. For $\lambda$, equation 5 expresses the boundary conditions that should be satisfied, where, $\lambda$ is the allowed wavelength of the field between the plates, $d$ is the distance between the plates, and $n$ is a positive integer ($n=1,2,3,$...) representing the mode number.
@@ 2.1.3 Repulsive Casimir effect
Although the Casimir force is attractive at the origin, a repulsive Casimir force was measured in (munday_measured_2009, ) when solid materials were immersed in a fluid. Munday, Capasso and Parsegian considered the dielectric response functions, $\epsilon$, of the materials. They observed that in a configuration of $\epsilon_{1}>\epsilon_{3}>\epsilon_{2}$ , where $\epsilon_{3}$ represents the medium and $\epsilon_{1}$ and $\epsilon_{2}$ the materials, the force is repulsive. These findings suggest applications such as quantum levitation, allowing the development of ultra-low friction devices and sensors (munday_measured_2009, ). In addition, the repulsive Casimir force was observed in (zhao_stable_2019, ) in a system consisting of a gold surface and a Teflon coated gold surface, immersed in ethanol.
Other methods for obtaining a repulsive Casimir force involve modifications of the geometry, such as the proposed glide-symmetric geometry (rodriguez_repulsive_2008, ). We observed that a change in geometry altered the allowed field modes. The repulsive Casimir force case is usually achieved by introducing an asymmetry between the two surfaces. Thus, both attractive and repulsive Casimir forces can be applied to MEMS.
@@ 2.1.4 Experiments and Materials
With a conductivity value of $4.5\times 10^{7}S/m$, gold is the most commonly used material because of its ability to be modified to achieve minimal roughness. Methods involve template stripping as suggested in (ederth_template-stripped_2000, ), achieving a root mean square roughness of less than $0.4$ nm, which is crucial for the accuracy of measurements at nanoscale distances. Moreover, (banishev_modifying_2012, ) distinguished the difference in the observed force when both bodies were made of Au, compared to when one body was made of Au and the other of Si. The force between the Au sphere and Si plate geometry with respect to the Au sphere and Au plate configuration was represented in a ratio of $0.74$ at separations of $100$ nm and $0.63$ at separations of $200$ nm , as recognized in (chen_investigation_2005, ).
Silver has an even higher conductivity value of $6.3\times 10^{7}S/m$, which should theoretically increase the magnitude of the Casimir force. However, owing to overtime oxidation, the surface roughness increases and, at the nanoscale, this will affect the accuracy of measurements, making gold preferred in experimental setups. The oxidation rate of silver increases with a decrease in the size of silver particles, notably for diameters $d_{p}\leq 6$ nm (chaparro_oxidation_2023, ). Furthermore, the oxidation process is shape-dependent, as described in (chaparro_shape-dependent_2024, ), where the oxidation process of silver nanostructures is observed, specifically for spheres, cubes, disks, cylinders, triangles, and pyramids. The nanospheres had the lowest level of oxidation among all the geometries tested (chaparro_shape-dependent_2024, ). Selecting specific silver nanostructures according to their oxidation characteristics can improve the design of MEMS and NEMS. Thus, the material properties affect the accuracy of the theoretical results compared to the experimental evaluation.
The Casimir force is observed as dominant in systems at the microscale and has been further experimentally verified. The Casimir force is observed as dominant in systems at the microscale and has been further experimentally verified. Lamoreaux conducted the first experimental measurement of the Casimir force between a flat plate and a sphere using a torsion pendulum for separations of $0.6$ to $6\mu m$. Both components of the composed geometry were coated with a $0.5\mu m$ Cu layer on each side, and further $0.5\mu m$ Au was evaporated on the faces undergoing attraction. The theoretical results agreed with the experimental measurements within an error of $5\%$ (lamoreaux_demonstration_1997, ). Another common method involves atomic force microscopy (AFM), owing to the high accuracy of measurements that it allows. U. Mohideen and A. Roy first implemented the AFM method for a plate-to-sphere geometry with a separation range of $0.1$ to $0.9\mu m$, and obtained an accuracy of $1\%$ (mohideen_precision_1998, ).
@@ 2.2 Theoretical Framework
First, a Casimir geometry consisting of an arc and a plate was introduced. The arc is essentially a finite curved nanomembrane, with length which we approximate to $6\mu m$ (due to the large radius of $R=100\mu m),$ concave to a finite plate with length 6 $\mu m$ (figure 1(a)). Both surfaces were fixed so that the Casimir interaction could not bring them together. The nanomembrane is modeled as a thin elastic arc with clamped endpoints, as a representation of a membrane fixed to a rigid MEMS frame. Clamped boundary conditions are assumed at the ends of the arc, which constrains both displacement and rotation. These assumptions are consistent with typical MEMS design (klimchitskaya_casimir_2024, ). Future work could extend these results through formal stability analysis. Figure 1(a) represents the arc-plate geometry at a separation of $0.1\mu m$ of the two surfaces, and figure 1(b) shows the arc-plate geometry at a distance $1\mu m$. Reference (gies_casimir_2006, ) shows that Casimir edge effects for finite plates are negligible when the finite extent of the plate in the direction parallel to its surface is sufficiently small. In this study, we consider the plate and arc to be sufficiently small in the direction parallel to their surfaces to satisfy this condition, allowing edge effects to be safely neglected throughout our analysis.
@@
The distance range between the arc and the plate, considered from the farthest point of the arc, is described as $d\in[0.1-1]\mu m$. Although the Casimir force increases with decreasing distance, most experiments operate in the micro- or sub-micro range, as in (mohideen_precision_1998, ), (bimonte_measurement_2021, ). Decreasing the distance to the nanoscale introduces challenges such as effects due to surface roughness and inability to achieve precision as the Casimir force becomes uncontrollably strong.
@@
The material choice is explained in Section 2.1.4. "Experiments and Materials" of this paper. The values for the Poisson’s ratio and Young’s modulus were taken into account in the case of a thin film for both gold and silver.